Diffusion Morning Report 03/10/05

Ted Berg tberg at snf.stanford.edu
Thu Mar 10 06:38:50 PST 2005


Good Morning All,
    Looking good this morning. SiGe has temp Stabilization issues 
-looking into it. Germane for Epi is low at 20 Psi-maybe we should 
change today?
Poly is at 27+ and will probably go down next week for pull and clean. 
Ebeam back in high vac.
   


  	Date: 	3/10/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	yellow 	Temp Stab Problem 	2.58/30
Tylan 9-Poly 	Up 	  	27.35/30
Tylan 10-Nitride 	Up 	  	24.2/30
Tylan 11-BPSG-LTO 	Up 	  	5.58/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	5.137/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well  	B on line 700 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050310/c50f0f1e/attachment.html>


More information about the diffusion mailing list